# Diff-BBO: Diffusion-Based Inverse Modeling for Black-Box Optimization

### v1

@@ Appendix B Analysis of Sub-optimality for Black-box Function
$\mathrm{Var}\left(\bm{x}_{k}^{*}-\bm{x}_{k}\right)$ $=\mathrm{Var}\left(\bm{x}_{k}^{*}\right)+\mathrm{Var}\left(\bm{x}_{k}\right)+\left({\mathbb{E}}[\bm{x}_{k}^{*}]-{\mathbb{E}}[\bm{x}_{k}]\right)^{2}-({\mathbb{E}}[\bm{x}_{k}^{*}-\bm{x}_{k}])^{2}$
$\leq\mathrm{Var}\left(\bm{x}_{k}^{*}\right)+\mathrm{Var}\left(\bm{x}_{k}\right)=2\mathrm{Var}\left(\bm{x}_{k}\right).$ (22)
By definition of $\Delta_{\mathrm{episdemic}}(\cdot)$, combining Equation 21 and Equation 21 completes the proof. Finally, we remark that for $N\geq 1$, $\bm{x}_{k}$ will be replaced by ${\mathbb{E}}{[\bm{x}_{j}]},\forall j\in[N]$, and is also consistent with the definition of $\Delta_{\mathrm{episdemic}}(\cdot)$. As such, the same conclusion holds for general $N$. ∎
@@ Appendix C Optimality of Proposed Acquisition Function
$\Leftrightarrow\max_{y_{k}\in\mathbb{R}}\sum_{k=1}^{K}y_{k}-\Delta(p_{\theta},y_{k}).$
By Section 4.4, which shows $\Delta(p_{\theta},y_{k}^{*})$ can be effectively upper bounded the epidemic uncertainty, we therefore have
$\max_{y_{k}\in\mathbb{R}}\sum_{k=1}^{K}f(\bm{x}_{k}),~{}~{}\bm{x}_{k}\sim p_{\theta}(\cdot\mid y_{k}),~{}~{}\theta\in\Theta$ $\Leftrightarrow\max_{y_{k}\in\mathcal{Y}}\sum_{k=1}^{K}y_{k}-\Delta_{\mathrm{episdemic}}(y_{k},\mathcal{D})$
Essentially, our chosen acquisition function allows Diff-BBO to maximize the lower bound of the original optimization problem. ∎
@@ Appendix D Experiment Details
DesignBench [Trabucco et al., 2022]
is a benchmark for real-world black-box optimization tasks. For continuouse tasks, we use Superconductor, D’Kitty Morphology and Ant Morphology benchmarks. For discrete tasks, we utilize TFBind8 and TFBind10 benchmarks. We exclude Hopper due to the domain is known to be buggy, as explained in Appendix C in [Krishnamoorthy et al., 2023]. We also exclude NAS due to the significant computational resource requirement. Additionally, we exclude the ChEMBL task because the oracle model exhibits non-trivial discrepancies when queried with the same design.
@@
We train our model on NVIDIA A100 GPU and report the average performance over 3 random runs along with standard deviation for each task. For discrete tasks, we follow the procedure in Krishnamoorthy et al. [2023] where we convert the $d$-dimensional vector to a $d\times c$ one hot vector regarding $c$ classes. We then approximate logits by interpolating between a uniform distribution and the one hot distribution using a mixing factor of $0.6$. We jointly train a conditional and unconditional model with the same model by randomly set the conditioning value to 0 with dropout probability of $0.15$.
@@
Appendix E Impact Statement
Optimization techniques can address various real-world problems, including drug and material design. Our method enhances sample-efficient online black-box optimization, potentially accelerating solutions in these areas. However, caution is needed to prevent misuse, such as optimizing drugs to enhance harmful side effects.

### v3

@@ Appendix B Analysis of Sub-optimality for Black-box Function
$\mathrm{Var}_{\theta,\theta^{*}}({\mathbb{E}}_{\bm{x}_{k},\bm{x}^{*}}[|f(\bm{x}^{*})-f(\bm{x}_{k})|~{}\rvert~{}\theta,\theta^{*}])$ $\leq\mathrm{Var}_{\theta,\theta^{*}}\left({\mathbb{E}}_{\theta,\theta^{*}}\left[8\sqrt{d}\sigma+\left\|{\mathbb{E}}_{\bm{x}^{*}}[\bm{x}^{*}]-{\mathbb{E}}_{\bm{x}_{k}}[\bm{x}_{k}]\right\|\right]\right)$
$\leq\mathrm{Var}_{\theta,\theta^{*}}\left(\left\|{\mathbb{E}}_{\bm{x}^{*}}[\bm{x}^{*}]-{\mathbb{E}}_{\bm{x}_{k}}[\bm{x}_{k}]\right\|\right).$
Then by property of variance, we have
$\mathrm{Var}_{\theta,\theta^{*}}\left(\left\|{\mathbb{E}}_{\bm{x}^{*}}[\bm{x}^{*}]-{\mathbb{E}}_{\bm{x}_{k}}[\bm{x}_{k}]\right\|\right)$ $={\mathbb{E}}_{\theta,\theta^{*}}\left[\left\|{\mathbb{E}}_{\bm{x}^{*}}[\bm{x}^{*}]-{\mathbb{E}}_{\bm{x}_{k}}[\bm{x}_{k}]\right\|^{2}\right]-\Big{(}{\mathbb{E}}_{\theta,\theta^{*}}\Big{[}\left\|{\mathbb{E}}_{\bm{x}^{*}}[\bm{x}^{*}]-{\mathbb{E}}_{\bm{x}_{k}}[\bm{x}_{k}]\right\|\Big{]}\Big{)}^{2}.$
From the proof of Lemma B.2, we have
$\quad{\mathbb{E}}_{\theta,\theta^{*}}\Big{[}\left\|{\mathbb{E}}_{\bm{x}^{*}}[\bm{x}^{*}|\theta^{*}]-{\mathbb{E}}_{\bm{x}_{k}}[\bm{x}_{k}|\theta]\right\|^{2}\Big{]}$
$={\mathbb{E}}_{\theta^{*}}[{\mathbb{E}}_{\bm{x}^{*}}[\left\|\bm{x}^{*}\right\|^{2}|\theta^{*}]]+{\mathbb{E}}_{\theta}[{\mathbb{E}}_{\bm{x}_{k}}[\left\|\bm{x}_{k}\right\|^{2}|\theta]]-2{\mathbb{E}}_{\theta,\theta^{*}}[{\mathbb{E}}_{\bm{x}_{k}}[(\bm{x}_{k}|\theta)]^{\mathrm{T}}{\mathbb{E}}_{\bm{x}^{*}}[\bm{x}^{*}|\theta^{*}]]$
$=2({\mathbb{E}}_{\theta}[{\mathbb{E}}_{\bm{x}_{k}}[\left\|\bm{x}_{k}\right\|^{2}|\theta]]-{\mathbb{E}}_{\theta,\theta^{*}}[{\mathbb{E}}_{\bm{x}_{k}}[(\bm{x}_{k}|\theta)]^{\mathrm{T}}{\mathbb{E}}_{\bm{x}^{*}}[\bm{x}^{*}|\theta^{*}]])$
$=2({\mathbb{E}}_{\theta}[{\mathbb{E}}_{\bm{x}_{k}}[\left\|\bm{x}_{k}\right\|^{2}|\theta]]-\left\|{\mathbb{E}}_{\theta}[{\mathbb{E}}_{\bm{x}_{k}}[\bm{x}_{k}|\theta]]\right\|^{2})$
$=2\mathrm{Var}_{\theta}({\mathbb{E}}_{\bm{x}_{k}}[\left\|\bm{x}_{k}]\right\|),$
where the third equality is by the law of total expectation and the fact that ${\mathbb{E}}_{\theta}[{\mathbb{E}}_{\bm{x}_{k}}[\bm{x}_{k}|\theta]]={\mathbb{E}}_{\theta^{*}}[{\mathbb{E}}_{\bm{x}^{*}}[\bm{x}^{*}|\theta^{*}]]$. Combining the above results, we have
$T_{2}\leq L^{2}\mathrm{Var}_{\theta,\theta^{*}}\left(\left\|{\mathbb{E}}_{\bm{x}^{*}}[\bm{x}^{*}]-{\mathbb{E}}_{\bm{x}_{k}}[\bm{x}_{k}]\right\|\right)\leq 2L^{2}\mathrm{Var}_{\theta}({\mathbb{E}}_{\bm{x}_{k}}[\left\|\bm{x}_{k}]\right\|).$ (29)
Combining Equation 28 and Equation 29 completes the proof:
$\mathrm{Var}\left(\Delta(p_{\theta},y_{k}^{*})\right)\leq c_{3}L^{2}d\sigma^{2}+2L^{2}\text{Var}_{\theta}({\mathbb{E}}_{\bm{x}}[\left\|\bm{x}_{k}\right\|]).$
B.1 Supporting Lemmas
Lemma B.3 (Wainwright [2019]).
Let $\bm{x}\in\mathbb{R}^{d}$ be a $\sigma$-subGaussian ramdom vector, then
${\mathbb{E}}[\left\|\bm{x}-{\mathbb{E}}[\bm{x}]\right\|]\leq 4\sigma\sqrt{d}.$ (30)
Lemma B.4.
Let $\bm{x}\in\mathbb{R}^{d}$ be a $\sigma$-subGaussian ramdom vector, then its variance satisfies:
$\text{Var}[\bm{x}]\leq Cd\sigma^{2},$ (31)
where C is some positive constant.
Proof of Lemma B.4.
By definition of sub-Gaussian vector, for any direction $\bm{u}\in\mathbb{R}^{d}$ with $\left\|\bm{u}\right\|=1$,
${\mathbb{E}}\left[\exp(\lambda\bm{u}^{\mathrm{T}}(\bm{x}-{\mathbb{E}}[\bm{x}]))\right]\leq\exp\left(\frac{\lambda^{2}\sigma^{2}}{2}\right),~{}~{}\forall\lambda\in\mathbb{R}.$
This implies that the second moment in any direction $\bm{u}$ satisfies:
${\mathbb{E}}\left[\bm{u}^{\mathrm{T}}((\bm{x}-{\mathbb{E}}[\bm{x}])(\bm{x}-{\mathbb{E}}[\bm{x}])^{\mathrm{T}})\right]\leq\sigma^{2}.$
Therefore, the maximum eigenvalue of the covariance matrix is upper-bounded by $C\sigma^{2}$, where $C$ is some positive constant.
$\text{Var}[\bm{x}]=\text{tr}\left({\mathbb{E}}\left[(\bm{x}-{\mathbb{E}}[\bm{x}])(\bm{x}-{\mathbb{E}}[\bm{x}])^{\mathrm{T}}\right]\right)\leq Cd\sigma^{2}.$
Lemma B.5.
In each iteration $k\in[K]$, let $\mathcal{D}$ be the collected dataset, $\theta$ and $\theta^{*}$ are parameters independently drawn from posterior $p(\theta|\mathcal{D})$, $\bm{x}_{k}\sim p_{\theta}(\cdot|y_{k}^{*},\mathcal{D})$ and $\bm{x}^{*}\sim p_{\theta^{*}}(\cdot|y_{k}^{*},\mathcal{D})$. For any measurable function $f$, and $\sigma(\mathcal{D})$-measurable random variable $\bm{x_{k}}$,
${\mathbb{E}}\left[f(\bm{x}_{k})\right]$ $={\mathbb{E}}\left[f(\bm{x}^{*})\right].$
Proof of Lemma B.5.
Since the black-box function $f$ is measurable, and by the nature of Algorithm 1, in each iteration $k$, the generated sample $\bm{x_{k}}$, the target function value $y_{k}^{*}$, the predictive distribution $p_{\theta}(\cdot|y_{k}^{*},\mathcal{D})$, the posterior distribution $p(\theta\mid\mathcal{D})$ are $\sigma(\mathcal{D})$-measurable at iteration $k$, the only randomness in $f(\bm{x})$ comes from the random sampling in the algorithm. Thus, condition on the training data $\mathcal{D}$ and target value $y_{k}^{*}$, by tower rule,
${\mathbb{E}}\left[f(\bm{x}_{k})\right]={\mathbb{E}}\left[{\mathbb{E}}\left[f(\bm{x}_{k})|\theta\right]\right]$ $=\int_{\theta}\int_{\bm{x}_{k}}f(\bm{x}_{k})p_{\theta}(\bm{x}_{k}|y_{k}^{*},\mathcal{D})p(\theta|\mathcal{D})\mathop{}\!\mathrm{d}\bm{x}_{k}\mathop{}\!\mathrm{d}\theta$
$=\int_{\theta}\int_{\bm{x}_{k}}f(\bm{x}_{k})p_{\theta}(\bm{x}_{k}|y_{k}^{*},\mathcal{D})\mathop{}\!\mathrm{d}\bm{x}_{k}~{}p(\theta|\mathcal{D})\mathop{}\!\mathrm{d}\theta.$
Note that both the true parameter $\theta^{*}$ and the chosen parameter $\theta$ are drawn from the same posterior distribution $p(\theta\mid\mathcal{D})$, we have
$\int_{\theta}\int_{\bm{x}}f(\bm{x})p_{\theta}(\bm{x}|y_{k}^{*},\mathcal{D})\mathop{}\!\mathrm{d}\bm{x}~{}p(\theta|\mathcal{D})\mathop{}\!\mathrm{d}\theta=\int_{\theta^{*}}\int_{\bm{x}}f(\bm{x})p_{\theta^{*}}(\bm{x}|y_{k}^{*},\mathcal{D})\mathop{}\!\mathrm{d}\bm{x}~{}p({\theta^{*}}|\mathcal{D})\mathop{}\!\mathrm{d}{\theta^{*}}.$
As a result, we have
${\mathbb{E}}\left[f(\bm{x}_{k})\right]$ $=\int_{\theta^{*}}\int_{\bm{x}^{*}}f(\bm{x}^{*})p_{\theta^{*}}(\bm{x}^{*}|y_{k}^{*},\mathcal{D})\mathop{}\!\mathrm{d}\bm{x}^{*}p({\theta^{*}}|\mathcal{D})\mathop{}\!\mathrm{d}{\theta^{*}}={\mathbb{E}}\left[{\mathbb{E}}\left[f(\bm{x}^{*})|\theta^{*}\right]\right]={\mathbb{E}}\left[f(\bm{x}^{*})\right].$
Corollary 1.
In each iteration $k\in[K]$, let $\mathcal{D}$ be the collected dataset, $\theta$ and $\theta^{*}$ are parameters independently drawn from posterior $p(\theta|\mathcal{D})$, $\bm{x}_{k}\sim p_{\theta}(\cdot|y_{k}^{*},\mathcal{D})$ and $\bm{x}^{*}\sim p_{\theta^{*}}(\cdot|y_{k}^{*},\mathcal{D})$. For any measurable function $f$, and $\sigma(\mathcal{D})$-measurable random variable $\bm{x_{k}}$,
${\mathbb{E}}\left[\left\|\bm{x}_{k}\right\|\right]$ $={\mathbb{E}}\left[\left\|\bm{x}^{*}\right\|\right].$
Proof of Corollary 1.
Since the norm function is deterministic and $\sigma(\mathcal{D})$-measurable, the proof directly follows that of Lemma B.5. ∎
@@ Appendix C Optimality of Proposed Acquisition Function
$\Rightarrow\max_{y_{k}\in\mathcal{Y}}\sum_{k=1}^{K}y_{k}-\Delta(p_{\theta},y_{k}),$
where the candidate set $\mathcal{Y}$ is constructed based on the training data and is designed to explore the objective space efficiently. When considering an online maximization problem, adding a positive term would lead to overestimation, because the model would be overly optimistic about $f(\bm{x}_{k})$. Therefore, we should only consider the case where we the uncertainty is being subtracted. By Section 5.4, which shows $\Delta(p_{\theta},y_{k}^{*})$ can be effectively upper bounded the epidemic uncertainty, we therefore have
$\max_{y_{k}\in\mathbb{R}}\sum_{k=1}^{K}f(\bm{x}_{k}),~{}~{}\bm{x}_{k}\sim p_{\theta}(\cdot\mid y_{k}),~{}~{}\theta\in\Theta$ $\Rightarrow\max_{y_{k}\in\mathcal{Y}}\sum_{k=1}^{K}y_{k}-\Delta_{\mathrm{episdemic}}(y_{k},\mathcal{D}).$
Essentially, our chosen acquisition function allows Diffusion-BBO to maximize the lower bound of the original optimization problem. Penalizing high uncertainty ensures that the model prioritizes more confident predictions (i.e. those with lower epistemic uncertainty), which are more likely to yield higher objective function values. ∎
@@ Appendix D Experiment Details
DesignBench [Trabucco et al., 2022] is a benchmark for real-world black-box optimization tasks. For continuouse tasks, we use Superconductor, D’Kitty Morphology and Ant Morphology benchmarks. For discrete tasks, we utilize TFBind8 and TFBind10 benchmarks. We exclude Hopper due to the domain is known to be buggy, as explained in Appendix C in [Krishnamoorthy et al., 2023]. We also exclude NAS due to the significant computational resource requirement. Additionally, we exclude the ChEMBL task because the oracle model exhibits non-trivial discrepancies when queried with the same design.
@@
Figure 4: Ablation study to evaluate the effect of batch size on the superconductor task. The mean and standard deviation across three random seeds are plotted. Diffusion-BBO shows robust performances across different batch size given the same total number of evaluations.
@@
We train our model on one NVIDIA A100 GPU and report the average performance over 3 random runs along with standard deviation for each task. For discrete tasks, we follow the procedure in Krishnamoorthy et al. [2023] where we convert the $d$-dimensional vector to a $d\times c$ one hot vector regarding $c$ classes. We then approximate logits by interpolating between a uniform distribution and the one hot distribution using a mixing factor of $0.6$. We jointly train a conditional and unconditional model with the same model by randomly set the conditioning value to 0 with dropout probability of $0.15$.
@@
Appendix E Ablation Study
We evaluate the effect of batch size, i.e., the number of queries per iteration on Diffusion-BBO on the Superconductor task. As shown in Figure 4, we compare the objective function score over number of function evaluations. We can see the performance of our approach remains similar when the batch size becomes larger, suggesting remarkable robustness across different batch sizes. Hence, Diffusion-BBO is a highly-scalable inverse modeling approach that can efficiently leverage parallelism to handle larger computational loads without compromising performance.